# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v3

@@ Abstract.
Scaling Large Language Model (LLM) training relies on multi-dimensional parallelism, where High-Bandwidth Domains (HBDs) are critical for communication-intensive parallelism like Tensor Parallelism (TP) and Expert Parallelism (EP). However, existing HBD architectures face fundamental limitations in scalability, cost, and fault resiliency: switch-centric HBDs (e.g., NVL-72) incur prohibitive scaling costs, while GPU-centric HBDs (e.g., TPUv3/Dojo) suffer from severe fault propagation. Switch-GPU hybrid HBDs such as TPUv4 takes a middle-ground approach by leveraging Optical Circuit Switches, but the fault explosion radius remains large at the cube level (e.g., 64 TPUs).
We propose InfiniteHBD, a novel transceiver-centric HBD architecture that unifies connectivity and dynamic switching at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, InfiniteHBD achieves reconfigurable point-to-multipoint connectivity, allowing the topology to adapt into variable-size rings. This design provides: i) datacenter-wide scalability without cost explosion; ii) fault resilience by isolating failures to a single node, and iii) full bandwidth utilization for fault-free GPUs. Key innovations include a Silicon Photonic (SiPh) based low-cost OCS transceiver (OCSTrx), a reconfigurable k-hop ring topology co-designed with intra-/inter-node communication, and an HBD-DCN orchestration algorithm maximizing GPU utilization while minimizing cross-ToR datacenter network traffic. The evaluation demonstrates that InfiniteHBD achieves 31% of the cost of NVL-72, near-zero GPU waste ratio (over one order of magnitude lower than NVL-72 and TPUv4), near-zero cross-ToR traffic when node fault ratios under 7%, and improves Model FLOPs Utilization by 3.37x compared to NVIDIA DGX (8 GPUs per Node).
copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price: 11footnotetext: Guyue Liu, Hao Nie, and Huaiyu Meng contributed equally to this work and share the corresponding authorship.
@@ 1. Introduction
Large-scale Large Language Models (LLMs) training rely on various parallelism strategies (Smith et al., 2022; Rajbhandari et al., 2020), such as Tensor Parallelism (TP), Expert Parallelism (EP), Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP) and Sequence Parallelism (SP). These strategies communicate over two types of AI datacenter compute fabrics, each with distinct bandwidth requirements. First, Datacenter Networks (DCNs) provide hundreds of Gbps per GPU and primarily handle DP, PP, CP, and SP traffic, which has lower communication demands. Second, High-Bandwidth Domains (HBDs) offer Tbps-level bandwidth, which is crucial for communication-intensive TP and EP. Efficient HBD design can reduce communication overhead, thereby improving Model FLOPs Utilization (MFU)-a key performance metric for LLM training.
The community has made significant advancements in designing DCNs for LLM training (Wang et al., 2024; Qian et al., 2024; NVIDIA, 2021; Gangidi et al., 2024a). However, scaling HBD to optimize MFU in LLM training remains a challenging problem. Existing HBD architectures (NVIDIA, 2024b; Jouppi et al., 2023; Tesla, 2024; Jouppi et al., 2020; Trainium, 2024) take important steps but still suffer from fundamental limitations in scalability, cost, and fault resiliency.
@@
    Switch-centric HBDs, such as NVIDIA NVL-72 (NVIDIA, 2024b), build multilayer nonconvergent networks for HBD with switch chips. However, the switch fabric incurs superlinear cost growth as it scales, constraining the number of GPUs per HBD. This limitation prevents optimal large TP and EP and causes severe resource fragmentation when the size of TP/EP group increases. For instance, with 2 HBDs (32 GPUs each), 30 GPUs are wasted for TP-16 jobs if each HBD has a single GPU failure. This waste reduces to 14 GPUs if the two HBDs are combined into a 64-GPU unit.
@@
    GPU-centric HBDs, such as Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021), adopts low-cost GPU-to-GPU links to construct large-scale ring or mesh topologies, forwarding traffic directly through GPUs. However, these architectures suffer from a large fault explosion radius, where a single GPU failure degrades bandwidth for a group of adjacent GPUs, compromising the entire topology. For example, in SiP-Ring, one single GPU failure breaks the ring and reforms the topology a line.
@@
    Switch-GPU Hybrid HBDs, TPUv4 (Jouppi et al., 2023) alleviates the limitation of large fault explosion radius via OCS-based switch [^1] In this paper, ”OCS” specifically denotes optical circuit switching capability, while OCS-based switch denotes optical circuit switch.: each set of 64 TPUs is connected as a cube, with these cubes connected to multiple OCS-based switches to isolate faults within the cubes. However, it cannot resolve it fundamentally, as the fault explosion radius remains large at the cube level (64 TPUs).
In this paper, we take a first-principles approach to redesigning HBD for LLM training workloads. Through a top-down analysis of parallelism strategies (§2.3) for maximizing MFU, we show that increasing TP size yields the most significant MFU gains for both dense and sparse LLM models. For large dense models (Llama Team, AI @ Meta, 2024), the optimal TP size scales from 16 to 64 as the number of GPUs increases. For sparse MoE models (Sun et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024b), enlarging TP improves MFU more effectively than EP, particularly when considering the expert imbalance problem (Liu et al., 2023b).
@@
    Silicon Photonics based OCS transceiver (OCSTrx): To design a cost-effective low-power transceiver with OCS support, we leverage the current advances of Silicon Photonics (SiPh) technology. Compared to MEMS (Urata et al., 2022a; Yeow et al., 2001) technology which has been widly used to realize OCS, SiPh offers simpler structures, lower cost and power consumption. We build OCS with Mach-Zehnder interferometer (MZI) matrix (Wikipedia, 2024), taped out with $65nm$ CMOS processes. The chip size is less than $136.5mm^{2}$ while the chip power consumption is $3.2Watts$, which can be integrated into commercial QSFP-DD $800Gbps$ transceiver (MSA, 2024) with sub-1ms path reconfiguration latency.
@@
To the best of our knowledge, InfiniteHBD is the first HBD design capable of scaling to datacenter scale while avoiding cost explosion and increased failure-induced waste. We evaluated InfiniteHBD with the real 348-day fault trace from our 10K GPU cluster [^2] Details in Appendix §A. We have open-sourced this trace at https://gitlab.basemind.com/sys/infra/superpod/infinitehbd-trace. When executing TP32 jobs with the trace, it demonstrates 0.15% GPU waste ratio - 69x and 45x lower than NVL-72 (10.31%) and TPUv4 (6.76%). It achieves 3.24x and 1.59x cost reductions compared to NVIDIA NVL-72 and Google TPUv4 respectively. Through the orchestration algorithm, it maintains near-zero cross-ToR traffic under 7% node failure rates. Its dynamic ring formation capability enables 3.37x higher MFU than NVIDIA DGX systems (Platform, 2024) (8 GPUs/node).
@@ 2. Background and Motivation
In this section, we first introduces LLM training in AI datacenters (DCs) (§2.1). Then, we examine existing High-Bandwidth Domain (HBD) architectures and discuss their limitations (§2.2). Finally, we summarize key design principles of HBD for LLM training (§2.3).
@@ 2.1. LLM Training in AI DC
LLM training parallelism and communication. LLM training jobs employ various parallelism strategies to efficiently utilize GPUs distributed across AI DCs (Shoeybi et al., 2020; Rajbhandari et al., 2020). Based on communication loads, parallelism can be categorized into two types. The first type is communication-intensive parallelism which involves high communication load. Tensor Parallelism (TP) splits the model across multiple GPUs and synchronizes via AllReduce. The ring algorithm for AllReduce is theoretically optimal (Patarasuk and Yuan, 2009), making ring-based topologies ideal for TP. Expert Parallelism (EP), designed for Mixture of Experts (MoE) models (Sun et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024b), assigns experts to different GPUs and relies on AlltoAll communication, requiring topologies with high bisection bandwidth (e.g., Full-Mesh). In contrast, parallelism strategies such as Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP), and Sequence Parallelism (SP) introduce lower communication overhead, placing less demands on network performance.
@@
Faults and fault explosion radius. As revealed by current advances of AI DCs (Qian et al., 2024; Gangidi et al., 2024b), training jobs experience a variety of faults, such as GPU faults, optical transceiver faults, switch faults, and link faults. We quantify the fault impact using the fault explosion radius, defined as the number of GPUs degraded by a single fault event. The fault explosion radius varies depending on both the system architecture and the fault component. For example, if a switch fails, the bandwidth of all devices connected to it will degrade, illustrating the switch-level fault explosion radius.
@@ 2.2. Limitations of Existing HBDs
Switch-centric HBD. This type architecture leverages switch chips to interconnect GPUs, as shown in Figure 1(a). A prominent example is NVIDIA, which utilizes NVLink and NVLink Switch  (NVIDIA, 2024d, 2018), e.g. DGX H100 (Platform, 2024) with 8-GPU and GB200 NVL-36, NVL-72, and NVL-576 (NVIDIA, 2024b). These architectures offer high-performance any-to-any communication. However, switch-centric HBDs have several drawbacks: i) They require a large number of switch chips due to their limited per-chip throughput; ii) They are vulnerable to a switch-level fault explosion radius—when a switch chip fails, all connected nodes experience bandwidth degradation; iii) High interconnect costs constrain the scale of HBDs, leading to significant fragmentation when serving large models.
GPU-centric HBD. GPU-centric HBD architectures construct the HBD using direct GPU-to-GPU connections, eliminating the need for switch chips. As a result, cost scales linearly with HBD size. A representative example is SiP-Ring (Khani et al., 2021), shown in Figure 1(b), where GPUs are organized into fixed-size rings. However, this design imposes a strict limitation: the TP group size must remain fixed. To enable communication at dynamic scales and support a wider range of workloads, more complex topologies are adopted (e.g., Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and AWS Trainium  (Trainium, 2024) ), which support dynamic scaling by allowing jobs to execute on topology subsets of varying sizes. As shown in Figure 1(c), Dojo (Tesla, 2024) connects GPUs via mesh-like topologies and employ GPUs to forward traffic. While GPU-centric architectures mitigate cost explosion and can support various scales, they suffer from a large fault explosion radius. A single GPU failure can disrupt the entire HBD by altering its connectivity, degrading communication performance even for healthy GPUs—such as the yellow GPUs in Figure 1(c).
Switch-GPU Hybrid HBD. This architecture interconnects GPUs via a combination of direct GPU-to-GPU connections and switch links. A typical example is TPUv4 (Jouppi et al., 2023), which organizes TPUs into $4^{3}$ TPU cubes and connect them via centralized OCS-based switches (Figure 1(d)). TPUv4 scales up to 4,096 TPUs, with its expansion primarily limited by the port count of the OCS-based switch. Furthermore, it suffers from a cube-level fault explosion radius—a failure in any single TPU affects the entire 64-TPU cube, leading to significant performance degradation. Furthermore, OCS-based switches face challenges of high costs and manufacturing complexity, which undermines the cost-effectiveness of TPUv4. TPUv5p cluster (Google, 2024) is similar to TPUv4 but can scale out to 8,960 TPUs.
@@ 2.3. Key Attributes of An Ideal HBD
Our production training experience on a 1T MoE model in production brings the following insights into the pros and cons of TP and EP. On the one hand, EP is more communication-efficient than TP. Table 3 compares the communication volume of TP and EP. Clearly, EP is better if $k<n$, which is common (Liu et al., 2024) because existing models often choose small $k$ for higher computation sparsity. On the other hand, EP suffers from the well-known expert imbalance problem (Liu et al., 2023b), especially when the MoE routers use the no-token-left-behind algorithm (Liu et al., 2024; Gale et al., 2023; Du et al., 2022). This will result in non-equivalent number of tokens that each expert will receive, which hence causes straggler nodes that waste GPU cycles of other nodes. Table 4 shows the simulated result of training GPT-MoE with 1.1T parameters (details in Appendix §B) under different expert imbalance coefficients [^4] Calculated as $\frac{max-min}{max}$, where $max$ and $min$ represent the maximum and minimum tokens allocated to each expert respectively.. When $coef=0$, EP is better than TP due to smaller communication overhead. As $coef$ increases, the MFU drops because of the straggler issue.
Key findings. These experiments provide us two key findings for HBD design. First, larger HBD size is increasingly needed for rapidly scaling LLMs (i.e., more than 1T parameters). Second, with larger HBD enabled, using TP is more favorable than EP to train MoE, because TP shards the computation equally across GPUs and hence bypasses the expert imbalance problem.
These findings reveal two key design principles for HBD: i) HBD must inherently support large and adaptable TP sizes, which fundamentally requires the scalability of HBD architecture; ii) the HBD designs need to ensure the effective support for the Ring-AllReduce communication. Given the demonstrated efficiency of TP in MoE training, ensuring support for Ring-AllReduce support is sufficient for mainstream LLM training scenarios; iii) small fault explosion radius. Thus, we propose designing a large and adaptable HBD architecture tailored for ring-based TP communication to optimize LLM parallelism strategies.

### v4

@@ Abstract.
Scaling Large Language Model (LLM) training relies on multi-dimensional parallelism, where High-Bandwidth Domains (HBDs) are critical for communication-intensive parallelism like Tensor Parallelism (TP) and Expert Parallelism (EP). However, existing HBD architectures face fundamental limitations in scalability, cost, and fault resiliency: switch-centric HBDs (e.g., NVL-72) incur prohibitive scaling costs, while GPU-centric HBDs (e.g., TPUv3/Dojo) suffer from severe fault propagation. Switch-GPU hybrid HBDs such as TPUv4 take a middle-ground approach, but the fault explosion radius remains large at the cube level (e.g., 64 TPUs).
We propose InfiniteHBD, a novel transceiver-centric HBD architecture that unifies connectivity and dynamic switching at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, InfiniteHBD achieves reconfigurable point-to-multipoint connectivity, allowing the topology to adapt to variable-size rings. This design provides: i) datacenter-wide scalability without cost explosion; ii) fault resilience by isolating failures to a single node, and iii) full bandwidth utilization for fault-free GPUs. Key innovations include a Silicon Photonic (SiPh)-based low-cost OCS transceiver (OCSTrx), a reconfigurable k-hop ring topology co-designed with intra-/inter-node communication, and an HBD-DCN orchestration algorithm maximizing GPU utilization while minimizing cross-ToR datacenter network traffic. The evaluation demonstrates that InfiniteHBD achieves 31% of the cost of NVL-72, near-zero GPU waste ratio (over one order of magnitude lower than NVL-72 and TPUv4), near-zero cross-ToR traffic when node fault ratios are under 7%, and improves Model FLOPs Utilization by 3.37x compared to NVIDIA DGX (8 GPUs per Node).
High-Bandwidth Domain, Optical Circuit Switching, Large Language Model
journalyear: 2025copyright: acmlicensedconference: ACM SIGCOMM 2025 Conference; September 8–11, 2025; Coimbra, Portugalbooktitle: ACM SIGCOMM 2025 Conference (SIGCOMM ’25), September 8–11, 2025, Coimbra, Portugaldoi: 10.1145/3718958.3750468isbn: 979-8-4007-1524-2/2025/09price: ccs: Networks Network architecturesccs: Networks Network componentsccs: Hardware Networking hardwareccs: Networks Data center networksccs: Computing methodologies Machine learning22footnotetext: Guyue Liu, Hao Nie, and Huaiyu Meng contributed equally to this work and share the corresponding authorship.
@@ 1. Introduction
Large Language Models (LLMs) training relies on various parallelism strategies (Smith et al., 2022; Rajbhandari et al., 2020), such as Tensor Parallelism (TP), Expert Parallelism (EP), Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP) and Sequence Parallelism (SP). These strategies communicate over two types of AI datacenter compute fabrics, each with distinct bandwidth requirements. First, Datacenter Networks (DCNs) provide hundreds of Gbps per GPU and primarily handle DP, PP, CP, and SP traffic, which has lower communication demands. Second, High-Bandwidth Domains (HBDs) offer Tbps-level bandwidth, which is crucial for communication-intensive TP and EP. Efficient HBD design can reduce communication overhead, thereby improving Model FLOPs Utilization (MFU)-a key performance metric for LLM training.
The community has made significant advancements in designing DCNs for LLM training (Wang et al., 2024; Qian et al., 2024; NVIDIA, 2021; Gangidi et al., 2024). However, scaling HBD to optimize MFU in LLM training remains a challenging problem. Existing HBD architectures (NVIDIA, 2024b; Jouppi et al., 2023; Tesla, 2024; Jouppi et al., 2020; Trainium, 2024) take important steps but still suffer from fundamental limitations in scalability, cost, and fault resiliency.
@@
    Switch-centric HBDs, such as NVIDIA NVL-72 (NVIDIA, 2024b), build multilayer non-blocking networks for HBD with switch chips. However, the switch fabric incurs superlinear cost growth as it scales, constraining the number of GPUs per HBD. This limitation prevents optimal large TP and EP and causes severe resource fragmentation when the size of TP/EP group increases. For instance, with 2 HBDs (32 GPUs each), 30 GPUs are wasted for TP-16 jobs if each HBD has a single GPU failure. This waste reduces to 14 GPUs if the two HBDs are combined into a 64-GPU unit.
@@
    GPU-centric HBDs, such as Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021), adopt low-cost GPU-to-GPU links to construct large-scale ring or mesh topologies, forwarding traffic directly through GPUs. However, these architectures suffer from a large fault explosion radius, where a single GPU failure degrades bandwidth for a group of adjacent GPUs, compromising the entire topology. For example, in SiP-Ring, one single GPU failure breaks the ring and reforms the topology into a line.
@@
    Switch-GPU Hybrid HBDs, TPUv4 (Jouppi et al., 2023) alleviates the limitation of large fault explosion radius via OCS-based switches [^1] In this paper, ”OCS” specifically denotes optical circuit switching capability, while OCS-based switch denotes optical circuit switch.: each set of 64 TPUs is connected as a cube, with these cubes connected to multiple OCS-based switches to isolate faults within the cubes. However, it does not fundamentally resolve the issue, as the fault explosion radius remains large at the cube level (64 TPUs).
In this paper, we take a first-principles approach to redesigning HBD for LLM training workloads. Through a top-down analysis of parallelism strategies (§2.3) for maximizing MFU, we show that increasing TP size yields the most significant MFU gains for both dense and sparse LLM models. For large dense models (Grattafiori et al., 2024), the optimal TP size scales from 16 to 64 as the number of GPUs increases. For sparse MoE models (Sun et al., 2024; DeepSeek-AI et al., 2025; Jiang et al., 2024b), enlarging TP improves MFU more effectively than EP, particularly when considering the expert imbalance problem (Liu et al., 2023b).
@@
    Silicon Photonics based OCS transceiver (OCSTrx): To design a cost-effective low-power transceiver with OCS support, we leverage the current advances of Silicon Photonics (SiPh) technology. Compared to MEMS (Urata et al., 2022; Yeow et al., 2001) technology which has been widely used to realize OCS, SiPh offers simpler structures, lower cost and power consumption. We build OCS with Mach-Zehnder interferometer (MZI) matrix (Wikipedia, 2024), taped out with $65nm$ CMOS processes. The chip area is smaller than $136.5mm^{2}$ while the chip power consumption is $3.2Watts$, which can be integrated into commercial QSFP-DD $800Gbps$ transceiver (MSA, 2024) with 60-80 $\mu s$ reconfiguration latency. OCSTrx introduces zero additional bit error rate in most cases, with an average insertion loss of 3.3dB at room temperature.
@@
To the best of our knowledge, InfiniteHBD is the first HBD design capable of scaling to datacenter scale while avoiding cost explosion and increased failure-induced waste. We evaluated InfiniteHBD with the real 348-day fault trace from our 3K-GPU cluster [^2] Details in Appendix §A. We have open-sourced this trace at https://github.com/stepfun-ai/InfiniteHBD-Trace.. When executing TP32 jobs with the trace, it demonstrates 0.53% GPU waste ratio - 20x and 14x lower than NVL-72 (10.04%) and TPUv4 (7.56%). It achieves 3.24x and 1.59x cost reductions compared to NVIDIA NVL-72 and Google TPUv4 respectively. Through the orchestration algorithm, it maintains near-zero cross-ToR traffic under 7% node failure rates. Its dynamic ring formation capability enables 3.37x higher MFU than NVIDIA DGX systems (Platform, 2024) (8 GPUs/node).
This work does not raise any ethical issues.
@@ 2. Background and Motivation
In this section, we first introduce LLM training in AI datacenters (DCs) (§2.1). Then, we examine existing High-Bandwidth Domain (HBD) architectures and discuss their limitations (§2.2). Finally, we summarize key design principles of HBD for LLM training (§2.3).
@@ 2.1. LLM Training in AI DC
LLM training parallelism and communication. LLM training jobs employ various parallelism strategies to efficiently utilize GPUs distributed across AI DCs (Shoeybi et al., 2020; Rajbhandari et al., 2020). Based on communication loads, parallelism can be categorized into two types. The first type is communication-intensive parallelism which involves high communication load. Tensor Parallelism (TP) splits the model across multiple GPUs and synchronizes via AllReduce. The ring algorithm for AllReduce is theoretically optimal (Patarasuk and Yuan, 2009), making ring-based topologies ideal for TP. Expert Parallelism (EP), designed for Mixture of Experts (MoE) models (Sun et al., 2024; DeepSeek-AI et al., 2025; Jiang et al., 2024b), assigns experts to different GPUs and relies on AlltoAll communication, requiring topologies with high bisection bandwidth (e.g., Full-Mesh). In contrast, parallelism strategies such as Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP), and Sequence Parallelism (SP) introduce lower communication overhead, placing less demands on network performance.
@@
Faults and fault explosion radius. As revealed by current advances of AI DCs (Qian et al., 2024; Gangidi et al., 2024), training jobs experience a variety of faults, such as GPU faults, optical transceiver faults, switch faults, and link faults. We quantify the fault impact using the fault explosion radius, defined as the number of GPUs degraded by a single fault event. The fault explosion radius varies depending on both the system architecture and the fault component. For example, if a switch fails, the bandwidth of all devices connected to it will degrade, illustrating the switch-level fault explosion radius.
@@ 2.2. Limitations of Existing HBDs
Switch-centric HBD. This type of architecture leverages switch chips to interconnect GPUs, as shown in Figure 1(a). A prominent example is NVIDIA, which utilizes NVLink and NVLink Switch  (NVIDIA, 2024d, 2018), e.g., DGX H100 (Platform, 2024) with 8-GPU and GB200 NVL-36, NVL-72, and NVL-576 (NVIDIA, 2024b). These architectures offer high-performance any-to-any communication. However, switch-centric HBDs have several drawbacks: i) They require a large number of switch chips due to their limited per-chip throughput; ii) They are vulnerable to a switch-level fault explosion radius—when a switch chip fails, all connected nodes experience bandwidth degradation; iii) High interconnect costs constrain the scale of HBDs, leading to significant fragmentation when serving large models.
GPU-centric HBD. GPU-centric HBD architectures construct the HBD using direct GPU-to-GPU connections, eliminating the need for switch chips. As a result, cost scales linearly with HBD size. A representative example is SiP-Ring (Khani et al., 2021), shown in Figure 1(b), where GPUs are organized into fixed-size rings. However, this design imposes a strict limitation: the TP group size must remain fixed. To enable communication at dynamic scales and support a wider range of workloads, more complex topologies are adopted (e.g., Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and AWS Trainium  (Trainium, 2024) ), which support dynamic scaling by allowing jobs to execute on topology subsets of varying sizes. As shown in Figure 1(c), Dojo (Tesla, 2024) connects GPUs via mesh-like topologies and employs GPUs to forward traffic. While GPU-centric architectures mitigate cost explosion and can support various scales, they suffer from a large fault explosion radius. A single GPU failure can disrupt the entire HBD by altering its connectivity, degrading communication performance even for healthy GPUs—such as the yellow GPUs in Figure 1(c).
Switch-GPU Hybrid HBD. This architecture interconnects GPUs via a combination of direct GPU-to-GPU connections and switch links. A typical example is TPUv4 (Jouppi et al., 2023), which organizes TPUs into $4^{3}$ TPU cubes and connects them via centralized OCS-based switches (Figure 1(d)). TPUv4 scales up to 4,096 TPUs, with its expansion primarily limited by the port count of the OCS-based switch. Furthermore, it suffers from a cube-level fault explosion radius—a failure in any single TPU affects the entire 64-TPU cube, leading to significant performance degradation. Furthermore, OCS-based switches face challenges of high costs and manufacturing complexity, which undermines the cost-effectiveness of TPUv4. TPUv5p cluster (Google, 2024) is similar to TPUv4 but can scale out to 8,960 TPUs.
@@ 2.3. Key Attributes of An Ideal HBD
Our production training experience on a 1T MoE model in production brings the following insights into the pros and cons of TP and EP. On the one hand, EP is more communication-efficient than TP. Table 3 compares the communication volume of TP and EP. Clearly, EP is better if $k<n$, which is common (DeepSeek-AI et al., 2025) because existing models often choose small $k$ for higher computation sparsity. On the other hand, EP suffers from the well-known expert imbalance problem (Liu et al., 2023b), especially when the MoE routers use the no-token-left-behind algorithm (DeepSeek-AI et al., 2025; Gale et al., 2023; Du et al., 2022). This will result in a non-equivalent number of tokens that each expert will receive, which hence causes straggler nodes that waste GPU cycles of other nodes. Table 4 shows the simulated result of training GPT-MoE with 1.1T parameters (details in Appendix §B) under different expert imbalance coefficients [^4] Calculated as $\frac{max-min}{max}$, where $max$ and $min$ represent the maximum and minimum tokens allocated to each expert respectively.. When $coef=0$, EP is better than TP due to smaller communication overhead. As $coef$ increases, the MFU drops because of the straggler issue.
Key findings. These experiments provide us with two key findings for HBD design. First, larger HBD size is increasingly needed for rapidly scaling LLMs (i.e., more than 1T parameters). Second, with larger HBD enabled, using TP is more favorable than EP to train an MoE model, because TP shards the computation equally across GPUs and hence bypasses the expert imbalance problem.
These findings reveal two key design principles for HBD: i) HBD must inherently support large and adaptable TP sizes, which fundamentally requires the scalability of HBD architecture; ii) the HBD designs need to ensure the effective support for the Ring-AllReduce communication. Given the demonstrated efficiency of TP in MoE training, ensuring support for Ring-AllReduce is sufficient for mainstream LLM training scenarios; iii) small fault explosion radius. Thus, we propose designing a large and adaptable HBD architecture tailored for ring-based TP communication to optimize LLM parallelism strategies.